OpenStudy (anonymous):

Write the equation of the transformed graph of sine with period PI that has been shifted horizontally to the right 3 units, has an amplitude of 3/4, and has been flipped across the x-axis.

OpenStudy (anonymous):

can you tell me how you find period?

OpenStudy (anonymous):

is it y=-3/4 sin(2-3)?

OpenStudy (anonymous):

you need an x first off all

OpenStudy (anonymous):

and which is that?

OpenStudy (anonymous):

alright so you have y=-3/4sin(bx+c)+d so after b should be an x

OpenStudy (anonymous):

i think it was a typo but it should be y=-3/4sin(2x-3)

OpenStudy (anonymous):

so the equation is right?

OpenStudy (anonymous):

if thats a horizontal shift it'd be y=-3/4sin(2x-6)